OpenStudy (anonymous):

Hey everyone back to school can someone help with this problem (x+2)^-1/2 (2x+4)+(x+2)^3/2

OpenStudy (anonymous):

Are you supposed to simplify the problem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok look at the parentheses on the left. Can you pull out any common factors?

OpenStudy (anonymous):

oh oops I mean one in the middle.

OpenStudy (anonymous):

ok but what do I do with those darn exponents. LOL

OpenStudy (anonymous):

just a sec, I'm working it out. will explain as soon as I'm done.

OpenStudy (anonymous):

ok here goes: Pulling out the common 2, you get: 2(x+2)(x+2)^-1/2 + (x+2)^3/2 right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So now can you see any terms that can be pulled out of the addition?

OpenStudy (anonymous):

actually first: can you simplify the left side of the addition at all?

OpenStudy (anonymous):

the (x+2)^-1/2 goes to the bottom and the 2 can be factored from the (2x+4) leaving 2(x+2) over (x+2)^1/2

OpenStudy (anonymous):

notice that both have a base of (x+2). What does that mean about the exponents?

OpenStudy (anonymous):

yea I'm sorry not a clue. Ugh Christmas break ruined me

OpenStudy (anonymous):

When the bases are equal, 2 multiplied exponents can be added. for example: 2^2 * 2^3 = 2^5. In this case, the base is (x+2) and the exponents are 1 and -1/2. soo.....

OpenStudy (anonymous):

I am so sorry and feel like an idiot right now. I get that the exponents are multiplied

OpenStudy (anonymous):

exponents should be added. In this case 1-1/2 = 1/2. As a result: you get 2(x+2)^1/2 + (x+2)^3/2 right?

OpenStudy (anonymous):

ok oops thats what I meant. But didn't the (-) exponent go to the denominator

OpenStudy (anonymous):

you can do that too, but in this case it's simply easier to add up the exponents, eliminating the denominator altogether.

OpenStudy (anonymous):

ok then what do I do?

OpenStudy (anonymous):

Now you have a common term of (x+2)^1/2 that you can pull out of both terms in the addition. It's like the opposite of the distributive property.

OpenStudy (anonymous):

Doing that, you get (x+2)^1/2 times(2 + x + 2), which simplifies to (x+4)(x+2)^1/2. and you're done.

OpenStudy (anonymous):

so my final answer should be (x+2)^1/2 + (x+4)

OpenStudy (anonymous):

yup! :)

OpenStudy (anonymous):

oh wait no. They're multiplied, not added.

OpenStudy (anonymous):

ok thanks (feel like an idiot) lol

OpenStudy (anonymous):

no problem!